OpenStudy (tygrr321):

what is the value of 17 14 (6.6x10 )-(9.2x10 ) --------------------- 16 (4x10 )

OpenStudy (tygrr321):

@ybarrap plz help

OpenStudy (tygrr321):

@jim_thompson5910

OpenStudy (tygrr321):

@Daniellelovee

jimthompson5910 (jim_thompson5910):

So the expression is this? \[\Large \frac{\left(6.6 \times 10^{17}\right)-\left(9.2 \times 10^{14}\right)}{4 \times 10^{16}}\]

OpenStudy (tygrr321):

yep

jimthompson5910 (jim_thompson5910):

it turns out that \[\Large 9.2 \times 10^{14} = 0.0092 \times 10^{17}\] notice how I moved the decimal point 3 spots to the left. To counter balance that, I bumped the exponent up 3 (from 14 to 17)

jimthompson5910 (jim_thompson5910):

so \[\Large \frac{\left(6.6 \times 10^{17}\right)-\left(9.2 \times 10^{14}\right)}{4 \times 10^{16}}\] turns into \[\Large \frac{\left(6.6 \times 10^{17}\right)-\left(0.0092 \times 10^{17}\right)}{4 \times 10^{16}}\]

OpenStudy (tygrr321):

so...um...?

jimthompson5910 (jim_thompson5910):

are you able to combine \[\Large \left(6.6 \times 10^{17}\right)-\left(0.0092 \times 10^{17}\right)\] ???

OpenStudy (tygrr321):

could u put it like i did because i dont understand that

jimthompson5910 (jim_thompson5910):

do you see how I got \[\Large 9.2 \times 10^{14} = 0.0092 \times 10^{17}\]

OpenStudy (tygrr321):

kinda

jimthompson5910 (jim_thompson5910):

do you see how I moved the decimal point 2 spots to the left when I went from 9.2 to 0.0092

OpenStudy (tygrr321):